A simple pendulum with a string length of 2.45 m has a maximum acceleration of 20 cm/s2. Find its speed when the position is 4 cm from the center.
Answer #1

Angular speed of simple pendulum,

w = sqrt (g / L)

w = sqrt (9.8 / 2.45) = 2 rad/s^2

Maximum acceleration, amax = Aw^2

0.20 = A*(2)^2

A = 0.05 m

Speed when  the position x = 4 cm = 0.04 m,

v = w sqrt (A^2 - x^2)

v = 2 * sqrt (0.052 - 0.042)

v = 0.06 m/s

v = 6 cm/s